OpenStudy (anonymous):

Ernie went to the store and bought 33 apples. He bought three times as many red apples as he bought green, and he bought one-half as many yellow as he did red. How many of each type of apple did he buy? (Points : 1) 16 red; 9 green; 8 yellow 21 red; 7 green; 5 yellow 18 red; 6 green; 9 yellow 24 red; 6 green; 3 yellow

OpenStudy (chosenmatt):

24 red; 6 green; 3 yellow @CausticSyndicalist @Kit_Kat_Nat_<3 think so?

OpenStudy (anonymous):

C is the correct answer Call green apples x. Three times as many red apples, so that's 3x. Half as many yellow as red, so that's 3x/2 x 3x 3x/2 = 33 11x/2 = 33 11x = 66 x = 6. 6 green apples, 3*6 = 18 red apples, and 18/2 = 9 yellow apples.

OpenStudy (anonymous):

so its out of c and d ?

OpenStudy (chosenmatt):

wait yea it is c

OpenStudy (anonymous):

c

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

np

OpenStudy (anonymous):

Fiona has a jar filled with 150 coins made up of quarters and nickels. The total value of the coins is $25.10. If x represents the number of quarters in the jar, which equation can be used to find the number of quarters? (Points : 1) 0.25x + 0.05(x – 150) = 25.10 0.05x + 0.25(150 – x) = 25.10 0.05x + 0.25(x – 150) = 25.10 0.25x + 0.05(150 – x) = 25.10

OpenStudy (anonymous):

B. 0.25x 0.05(150 - x) = 26.70 Quarters = x, so nickels = 150 - x Quarters are 25 cents each, so they're worth 0.25x Nickels are 5 cents each, so they're worth 0.05(150 - x) 0.25x 0.05(150 - x) = 26.70

OpenStudy (anonymous):

did that help?

OpenStudy (anonymous):

where did you get 26.0 ?

OpenStudy (anonymous):

26.70*

OpenStudy (anonymous):

yeah it help

OpenStudy (anonymous):

Hannah has a jar filled with 150 coins made up of quarters and nickels. The total value of the coins is $28.30 How many quarters and how many nickels are in the jar? (Points : 1) 46 quarters; 104 nickels 75 quarters; 75 nickels 104 quarters; 46 nickels 95 quarters; 55 nickels

OpenStudy (anonymous):

104 quarters; 46 nickels